Ravenna Company is a merchandiser that uses the indirect method to prepare the operating activities section of its statement of cash flows. Its balance sheet for this year is as follows:

Ending Balance Beginning Balance
Cash and cash equivalents $ 102,000 $ 122,400
Accounts receivable 81,700 88,000
Inventory 109,700 100,000
Total current assets 293,400 310,400
Property, plant, and equipment 291,000 280,000
Less accumulated depreciation 97,000 70,000
Net property, plant, and equipment 194,000 210,000
Total assets $ 487,400 $ 520,400
Accounts payable $ 64,000 $ 113,700
Income taxes payable 49,700 65,700
Bonds payable 120,000 100,000
Common stock 140,000 120,000
Retained earnings 113,700 121,000
Total liabilities and stockholders’ equity $ 487,400 $ 520,400

During the year, Ravenna paid a $12,000 cash dividend and it sold a piece of equipment for $6,000 that had originally cost $13,800 and had accumulated depreciation of $9,200. The company did not retire any bonds or repurchase any of its own common stock during the year.

8-a. If the company debited income tax expense and credited income taxes payable $1,510 during the year, what is the total amount of the debits recorded in the Income Taxes Payable account?

8-b. What does the amount of these debits represent?

9-a. What is the amount and direction (+ or −) of the income taxes payable adjustment to net income in the operating activities section of the statement of cash flows?

9-b. What does this adjustment represent

10. Would the operating activities section of the company’s statement of cash flows contain an adjustment for a gain or a loss? What would be the amount and direction (+ or −) of the adjustment?
Answer #1

ANSWER

8-a) income tax payable
cash 5,920 opening 65,700
year 1,510
end 49,700
total debits recorded 17,510 answer
8-b) These debits represent cash paid to reduce the income
tax outstanding

9a)

Decrease in income tax payable = 65,700-49,700= -16,000

Adjustment to reconcile net income to net cash flow from operating activities
Decrease in income tax payable -16,000

9b)

So this adjustment represent cash outflow from Cash paid for income tax..

10)

Book value of equipment = 13,800 -9,200= 4,600

Gain on sale of equipment = Sale value-Book value = 6,000 -4,600= 1,400

Adjustment to reconcile net income to net cash flow from operating activities
Gain on sale of equipment -1,400

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************